# H3AE: High Compression, High Speed, and High Quality AutoEncoder for Video Diffusion Models

### v1

@@ 3.1.3 Efficient and High Compression Autoencoder
Wrapping up the design, we divide both the encoder and the decoder into two stages. The high-resolution stage utilizes spatial Conv2D for sliced inference, the spatial-temporal stage uses 3D Causal convolution blocks as well as 3D Causal Attention to model temporal dependency and capture global receptive field, as in Fig. 2. Note that even though multiple downsampling (upsampling) sub-stages are involved in the spatial-temporal stage, we only distribute 3D Causal Attention in the bottleneck position with the highest compression ratio. This straightforward design aligns the computation and memory complexity of the 3D Attentions in VAE with those in the DiT denoiser, ensuring that the entire diffusion pipeline is runnable on resource-constrained devices. For instance, our $8\times 32\times 32$ H3AE has only $1/32$ tokens in the bottleneck and latent stage compared to the CogVideoX $4\times 8\times 8$ VAE, thus the 3D Attentions in our H3AE and our DiT enjoy more than 1000×1000$\times$1000 × FLOPs reduction compared to CogVideoX DiT, significantly improving computational efficiency for video generation.
Besides architecture design, it is also crucial to properly distribute network parameters and computations to achieve a better Pareto curve of performance and efficiency. We investigate the depth and width scaling property of our VAE by directly profiling on the mobile device (iPhone 16 PM). We intuitively set the maximum network width based on the memory bound, and set a real-time target for mobile inference. i.e., decoding a $17\times 512\times 512$ video clip within 0.5s, which is equivalently more than 30 FPS. We scale the network by shrinking the width but increasing the depth to maintain this real-time target, and explore reconstruction performance, as in Tab. 2. We find that in this scope, wide but shallow VAE achieves the best reconstruction quality. All of our H3AEs are then constructed based on depth-width ratios of the 1×1$\times$1 × variant in Tab. 2.
@@
0.5×0.5$\times$0.5 × 233.0 0.45 27.02
0.75×0.75$\times$0.75 × 206.1 0.46 29.11
1×1$\times$1 × 196.5 0.45 29.48
Table 2: Scaling of network backbone. The 1×1$\times$1 × backbone is constructed under real-time constraint while maximizing the width to fit mobile memory.
@@
$8\times 32\times 32$ 29.48 0.8280 0.2627 0.2611 147.1
omni-plain 29.95 0.8367 0.2506 0.2470 129.6
omni-I2V 30.08 0.8384 0.2456 0.2427 128.4
$8\times 64\times 64$ 25.21 0.7284 0.3937 0.4048 737.9
omni-plain 25.97 0.7387 0.3963 0.4041 709.7
omni-I2V 26.38 0.7503 0.3748 0.3879 684.1
Table 3: Omni-AE for the T2V and I2V tasks. The baseline is trained without image condition while omni-training utilized the image condition with probability ($p=0.5$). The result shows that adding the omni-training objective helps both plain reconstruction and image-conditioned setting for I2V.
3.2 Omni-AE for both T2V and I2V
Image-to-video (I2V) generation is a prominent application in video generation [65, 71, 57], aiming to synthesize realistic and temporally coherent video sequences conditioned with a user-specified input image. This task is crucial for various applications, including video prediction, animation, content creation, and scene understanding.
In the I2V task, an image (typically the first frame) is provided as the generation condition. Tian et al. [57] propose to construct an image-conditioned VAE to utilize this additional information, achieving better decoding quality in high compression settings. Though utilizing the image condition benefits quality, the dedicated nature of I2V-VAE prohibits its wide deployment. Considering the cost of adapting the denoiser to a new VAE when swapping from T2V to I2V, most video diffusion works [5, 22, 46] still use plain VAEs for simplicity.
In this work, we propose a simple but effective multifunctional VAE that works for both plain T2V and conditioned I2V setting. Specifically, we propose an omni-objective training, where we take the hierarchical features of the first frame from the encoder and feed them to the decoder as condition with probability $p$. The decoder simulates I2V scenario when receiving the condition features, while performing plain reconstruction when not. As shown in Tab. 3, we find that after the omni-objective training, our VAE can perform both plain reconstruction and image-conditioned reconstruction effectively with a single set of weights. Interestingly, because of the auxiliary information, not only does the I2V setting get enhanced quality, but we also find that this training method improves plain reconstruction. This finding shows that it is a free lunch to train Omni-AEs serving both tasks: plain reconstruction and I2V setting. For the main experimental results in the following section, unless otherwise stated, we train H3AEs with this omni-objective and inference with the plain reconstruction setting to fairly compare with baselines.
[…]
Baseline 29.48 0.8280 0.2627 0.2611 147.1
LPIPS 29.07 0.8205 0.224 0.2275 148.9
GAN 29.35 0.8241 0.2631 0.2623 161.0
DWT 29.36 0.8247 0.2651 0.2642 159.3
Ours 29.58 0.8299 0.2541 0.2511 142.2
Table 4: Training loss comparison. Baseline is $L_{\text{recon}}+\lambda L_{\text{KL}}$. Ours represent the baseline followed by fine-tuning with the latent consistency loss. It shows that our proposed loss function yields better reconstruction quality (measured by rFVD) and achieves considerably better metrics than other auxiliary losses.
Model $f$Params. (M) FPS@512×512@$\times 512\uparrow$ 512 ↑ DAVIS OpenVid-HD
@@ 3.2 Omni-AE for both T2V and I2V
H3AE $8\times 64\times 64$ 643.76 182.8 39.0 26.03 0.7404 688.4 30.23 0.8788 39.3
Table 5: SOTA AE Comparison. Comparing with state-of-the-art autoencoders on latency, reconstruction quality on DAVIS and OpenVid datasets. For a fair comparison, we report our results without image condition (plain VAE). The I2V performance of our VAE is shown in Tab. 3 and Fig. 5. Note that our VAE is fully causal.
3.3 Training Losses
Existing works have investigated various auxiliary losses such as Perceptual loss [26], Discrete Wavelet Transform (DWT) loss [37, 18], GAN loss [18, 31, 65], etc. Perceptual loss measures the reconstruction error in feature space of a pre-trained neural network. Similarly, DWT loss computes the feature difference in a wavelet frequency space, while GAN loss introduces a discriminator to learn the separation between real and fake samples. In our empirical study on large-scale video training, we find that these auxiliary losses hardly provide improvements over the baseline, as shown in Tab. 4. Plus, most of them substantially slow down the training because of the large computation graph, and brings visual artifacts as also noticed by [18].
3.3.1 Latent Consistency Loss
To overcome the drawbacks of these discriminative losses, we design a new auxiliary training loss to enhance reconstruction quality by utilizing the unique paradigm of VAEs. Specifically, we reuse the VAE encoder as the discriminator, and encode the reconstructed video to obtain a fake posterior ($z^{{}^{\prime}}$). Note that the VAE encoder weights are not updated at this step. We compare the KL Divergence between the fake posterior ($z^{{}^{\prime}}$) and the posterior ($z$) encoded from the ground truth video, which has already been computed in the training forward pass. Thus, the latent consistency loss (LC loss) is:
$\begin{split}\mathcal{L_{\text{LC}}}=&D_{KL}(z,z^{\prime})\\ =&\frac{1}{2}\sum\left(\frac{(\mu_{z}-\mu_{z^{\prime}})^{2}}{\sigma_{z^{\prime}}^{2}}+\frac{\sigma_{z}^{2}}{\sigma_{z^{\prime}}^{2}}-1-\log\frac{\sigma_{z}^{2}}{\sigma_{z^{\prime}}^{2}}\right)\end{split}$ (2)
This design holds several advantages. First, we eliminate the need to design complicated discriminators as in GAN loss, and the training is more stable without the risk of mode collapse. Second, the discriminative loss provides stronger guidance compared to LPIPS and DWT, as the encoder is inherited on the fly from the VAE itself, which seamlessly preserves the knowledge and evolves itself. As in Tab. 4, we show that our proposed LC loss consistently improves reconstruction.
@@ 3.3.1 Latent Consistency Loss
Figure 3: AE Qualitative Results. Reconstructions from our H3AE ($8\times 32\times 32$) and other high compression autoencoders: Cosmos-Tokenizer [1] ($8\times 16\times 16$), LTX-VAE [18] ($8\times 32\times 32$). We show zoomed-in results to highlight the differences in fidelity and quality. Our method features greater high-frequency detail. Here ‘GT‘ refers to the ground truth video.
4 Experiments
@@
Table 6: Quantitative Metrics on VBench. Text-to-video generation benchmark on Vbench [24] with a 2B DiT denoiser trained using different Autoencoders. Overall Vbench score and selected score are reported. We specifically exhibit aesthetic quality (AQ), imaging quality (IQ), and motion smoothness (MS) to evaluate the performance of denoisers. Latency of DiT is benchmarked by generating 65-frame $512\times 512$ video clips on an NVIDIA A100 80GB GPU and 17-frame $512\times 512$ clips on iPhone 16 Pro Max. Notably the DiT using Cosmos-Tokenizer results in OOM error on iPhone due to memory inefficiency.
@@
Implementation details.  The model is trained using our internal video dataset. We resize and randomly crop a $256\times 256$ patch from each video clip for training. The model is initially trained using 17-frame video clips for 70K iterations for faster convergence, and then finetuned on video clips with various lengths (1-65) for 10K iterations for better generalization. Latent consistency loss is only applied in the second stage. The training is conducted on 32 NVIDIA A100 80G GPUs, with a batch size of 16 per GPU. We use AdamW optimizer with $41e-$ 4 learning rate and $\beta=[0.9,0.999]$ for the first stage, and reduce the learning rate to $51e-$ 5 in the second stage.
Evaluation.  We benchmark reconstruction quality using high resolution video dataset, including 60 video clips from DAVIS-2017 [44] testset and 4000 high resolution video clips randomly sampled from OpenVid [42] (OpenVid data is excluded in our training to ensure zero-shot evaluation). The first 33 frames from each clip are utilized for evaluation with a spatial resolution at $512\times 512$. We evaluate the PSNR, SSIM, and reconstruction-FVD (rFVD) to benchmark the performance of autoencoders.
4.1 Comparison with SOTA autoencoders
We compare our model with SOTA video tokenizers, i.e. CogVideoX-VAE, CV-VAE [69], Cosmos-Tokenizer [1], LTX-VAE [18], etc. Among them, Cosmos-Tokenizer and LTX-VAE are focused on high compression ratios. Three variants of H3AE models are built up with $4\times 16\times 16$, $8\times 32\times 32$, and $8\times 64\times 64$ compression ratios respectively to demonstrate the robustness of our design choices and analysis. We obtain the pre-trained SOTA models and evaluate them under the same evaluation setting mentioned above for reconstruction quality. We report inference speed on Nvidia A100 GPU and iPhone 16 Pro.
As shown in Tab. 5 and Fig. 1, our models achieve better reconstruction under the same or higher compression ratio. Specifically, with 4×4$\times$4 × higher compression ratio, our $4\times 16\times 16$ VAE outperforms Cosmos-Tokenizer ($4\times 8\times 8$) across both evaluation sets, achiving notable improvements of $+1.76$ PSNR, $+0.0333$ SSIM, $-14.22$ rFVD on DAVIS, as well as and $+2.27$ PSNR, $+0.0174$ SSIM, $-0.96$ rFVD on OpenVid-HD. Furthermore, our $8\times 32\times 32$ VAE significantly outperforms LTX-VAE by $+1.89$ PSNR, $+0.048$ SSIM, $-42.61$ rFVD on DAVIS, along with $+0.61$ PSNR, $+0.0152$ SSIM, $-0.27$ rFVD on OpenVid-HD. In addition, our model demonstrates superior overall efficiency. Our $8\times 32\times 32$ VAE requires only half the parameters compared to LTX-VAE, and achieves 2.5×2.5$\times$2.5 × speedup faster on iPhone, making it significantly more lightweight and faster for real-time applications.
4.2 Video Generation
We show text-to-video generation results in Tab. 6 and Fig. 4 to demonstrate that our VAE creates a good and highly-compressed latent space for video diffusion models. We construct a DiT with 2B parameters aligned with popular video diffusion model settings. Specifically, we use 3D full attention with RoPE and QK-Norm in transformer blocks. All models are first trained on image dataset to learn spatial knowledge for $50K$ iterations, then these model are trained with image-video joint training strategy for $100K$ iterations. The VDM training is done on 128 A100 GPUs for 4-10 days depending on the VAE compression ratio. As shown in Tab. 6, we find that using our H3AE ($8\times 32\times 32$) outperforms Cosmos-Tokenizer ($8\times 16\times 16$) and LTX-VAE in Vbench score. Our high-compression ratio autoencoder also delivers fast and memory efficient inference on both GPU and edge device. We additionally report the result of using an extremely high-compression ratio ($8\times 64\times 64$) H3AE, which provides more than a 4×4$\times$4 × speed-up on the iPhone 16 PM. This demonstrates the great potential of our approach for efficient video generation on edge devices.

### v2

@@ 3.1.2 Macro Architecture
Wrapping up the design, we divide both the encoder and the decoder into two stages. The high-resolution stage utilizes spatial Conv2D for sliced inference, the spatial-temporal stage uses 3D Causal convolution blocks as well as 3D Causal Attention to model temporal dependency and capture global receptive field, as in Fig. 2. Note that even though multiple downsampling (upsampling) sub-stages are involved in the spatial-temporal stage, we only distribute 3D Causal Attention in the bottleneck position with the highest compression ratio. This straightforward design aligns the computation and memory complexity of the 3D Attentions in VAE with those in the DiT denoiser, ensuring that the entire diffusion pipeline is runnable on resource-constrained devices. For instance, our $8\times 32\times 32$ H3AE has only $1/32$ tokens in the bottleneck and latent stage compared to the CogVideoX $4\times 8\times 8$ VAE, thus the 3D Attentions in our H3AE and potential DiT denoiser enjoy more than $1000\times$ FLOPs reduction compared to CogVideoX DiT, significantly improving computational efficiency for video generation.
Table 2: Scaling of network backbone. The $1\times$ backbone is constructed under real-time constraint while maximizing the width to fit mobile memory.
@@
$0.5\times$ 233.0 0.45 27.02
$0.75\times$ 206.1 0.46 29.11
$1\times$ 196.5 0.45 29.48
Besides architectural design, it is also crucial to properly distribute network parameters and computations to achieve a better Pareto curve of performance and efficiency. We investigate the depth and width scaling property of our VAE by directly profiling on the mobile device (iPhone 16 PM). We intuitively set the maximum network width based on the memory bound, and set a real-time target for mobile inference. i.e., decoding a $17\times 512\times 512$ video clip within 0.5s, which is equivalently more than 30 FPS. We scale the network by shrinking the width but increasing the depth to maintain this real-time target, and explore reconstruction performance, as in Tab. 2. We find that in this scope, a wide but shallow VAE achieves the best reconstruction quality. All of our H3AEs are then constructed based on depth-width ratios of the $1\times$ variant in Tab. 2.
3.2 Omni-AE for both T2V and I2V
Image-to-video (I2V) generation is a prominent application in video generation (Yang et al., 2024; Zheng et al., 2024; Tian et al., 2024), aiming to synthesize video sequences conditioned with a user-specified input image (typically the first frame). This task is crucial for various applications such as video prediction and animation. Tian et al. (Tian et al., 2024) propose to construct an image-conditioned VAE to utilize this additional information, achieving better reconstruction quality in high compression settings. Though utilizing the image condition benefits quality, the dedicated nature of I2V-VAE prohibits its wide deployment. Considering the cost of adapting the denoiser to a new VAE when swapping from T2V to I2V, most video diffusion works (Blattmann et al., 2023; Hong et al., 2022; Polyak et al., 2024) still use plain VAEs for simplicity.
In this work, we propose a simple yet effective multifunctional VAE that works for both plain T2V and conditioned I2V settings. Specifically for the decoder ($\hat{x}\leftarrow\mathrm{Decoder}(z)$), we propose an omni-objective training, where we take the hierarchical features ($e_{i}$) of the first frame from the encoder and feed them to the decoder as condition with probability $p$.
$\hat{x}=\begin{cases}\displaystyle\prod_{i}\mathrm{DecoderBlock}_{i}(z_{i}),&\text{with probability }1-p,\\ \displaystyle\prod_{i}\mathrm{DecoderBlock}_{i}\bigl((z_{i}+e_{-i})/2\bigr),&\text{with probability }p,\end{cases}$ (1)
where $z_{i}$ and $e_{-i}$ are decoder and encoder feature at symmetric positions. The decoder simulates I2V scenario when receiving the condition features, while performing plain reconstruction when not. As shown in Sec. 4.2, we find that after the omni-objective training, our VAE can perform both plain reconstruction and image-conditioned reconstruction effectively with a single set of weights. Interestingly, because of the auxiliary information, not only does the I2V setting get enhanced quality, but we also find that this training method improves plain reconstruction. This finding shows that it is a free lunch to train Omni-AEs serving both tasks: plain reconstruction and I2V setting. For the main experimental results in the following section, unless otherwise stated, we train H3AEs with this omni-objective and inference with the plain reconstruction setting to fairly compare with baselines.
3.3 Latent Consistency Loss
Existing works have investigated various auxiliary losses such as Perceptual loss (Johnson et al., 2016), Discrete Wavelet Transform (DWT) loss (Lin et al., 2024; HaCohen et al., 2024), GAN loss (HaCohen et al., 2024; Kong et al., 2024; Yang et al., 2024), etc. Perceptual loss measures the reconstruction error in feature space of a pre-trained neural network. Similarly, DWT loss computes the feature difference in a wavelet frequency space, while GAN loss introduces a discriminator to learn the separation between real and fake samples. In our empirical study on large-scale video VAE training, we find that the performance gain from these auxiliary losses are limited, as in Sec. 4.2. Plus, it is very likely to brings grid artifacts as in Fig. 3, and also noticed by HaCohen et al. (2024).
To overcome the drawbacks of these auxiliary losses, we design a new auxiliary training loss to enhance reconstruction quality by utilizing the unique paradigm of VAEs. Specifically, we reuse the VAE encoder as the discriminator, and encode the reconstructed video to obtain a fake posterior ($z^{{}^{\prime}}$). Note that the VAE encoder weights are not updated at this step. We compare the KL Divergence between the fake posterior ($z^{{}^{\prime}}$) and the posterior ($z$) encoded from the ground truth video, which has already been computed in the training forward pass. Thus, the latent consistency loss (LC loss) is:
$\mathcal{L_{\text{LC}}}=D_{KL}(z,z^{\prime})=\frac{1}{2}\sum\left(\frac{(\mu_{z}-\mu_{z^{\prime}})^{2}}{\sigma_{z^{\prime}}^{2}}+\frac{\sigma_{z}^{2}}{\sigma_{z^{\prime}}^{2}}-1-\log\frac{\sigma_{z}^{2}}{\sigma_{z^{\prime}}^{2}}\right)$ (3)
This design holds several advantages. First, we eliminate the need to incorporate an extra discriminative network as in GAN and LPIPS training. Second, empirically we found that the training is more stable without the risk of mode collapse or bringing in certain artifact patterns, as the qualitative visualizations. Lastly, latent consistency loss provides a stronger guidance as the discriminator is inherited on the fly from the VAE encoder which is more powerful than a frozen network (e.g., VGG for LPIPS loss). In Sec. 4.2, we show that our proposed latent consistency loss improves all reconstruction and fidelity metrics.
4 Experiments
Implementation details.  H3AE is trained on our internally collected image and video dataset, which has similar context statistics and aesthetics to public large-scale datasets such as Chen et al. (2024). Due to the absence of commonly adopted large-scale video dataset and different policies, most SOTA Video VAEs (Wan et al., 2025; HaCohen et al., 2024; Agarwal et al., 2025) are trained on their private dataset, making fully fair comparison difficult. To overcome this, we retrain LTX-VAE on our dataset with our training recipe, and obtain very close results, as in Appendix Tab. 8. We argue that due to the reconstruction nature, VAE training is less sensitive to video data quality and distribution. On the other hand, all design ablations in this work, including the architecture, omni-objective training and latent consistency loss are trained on the same data and recipe so fair comparison is ensured. Our model is trained on $256\times 256$ video clips with various length from 1 to 49 for 80K iterations. Latent consistency loss is only applied in the final 10K iters. The training is conducted on 32 NVIDIA A100 80G GPUs. We use AdamW optimizer with 1e41e-4 learning rate and $\beta=[0.9,0.999]$, and reduce the learning rate to 1e51e-5 in the last 10K iters.
Evaluation.  We benchmark reconstruction quality using high resolution video dataset, including 60 video clips from DAVIS-2017 (Perazzi et al., 2016) testset and 4000 high resolution video clips randomly sampled from OpenVid (Nan et al., 2025) Our training data does no overlap with DAVIS or OpenVid, so that zero-shot evaluation is strictly enforced. The first 33 frames from each clip are utilized for evaluation with a spatial resolution at $512\times 512$. We evaluate the PSNR, SSIM, and reconstruction-FVD (rFVD) to benchmark the performance of autoencoders.
Table 3: Comparison with SOTA Autoencoders. Comparing with state-of-the-art autoencoders on latency, reconstruction quality on DAVIS and OpenVid datasets. For a fair comparison, we report our results without image condition (plain VAE). The I2V performance of our VAE is shown in Sec. 4.2 and Fig. 5. Note that our VAE is fully causal.
@@ 4 Experiments
H3AE $8\times 64\times 64$ 643.76 182.8 39.0 26.03 0.7404 688.40 30.23 0.8788 39.3
4.1 Comparison with SOTA autoencoders
We compare our model with SOTA video tokenizers, i.e. CogVideoX-VAE, CV-VAE (Zhao et al., 2024a), Cosmos-Tokenizer (Agarwal et al., 2025), LTX-VAE (HaCohen et al., 2024), etc. Among them, Cosmos-Tokenizer and LTX-VAE are focused on high compression ratios. Three variants of H3AE models are built up with $4\times 16\times 16$, $8\times 32\times 32$, and $8\times 64\times 64$ compression ratios respectively to demonstrate the robustness of our design choices and analysis. We obtain the pre-trained SOTA models and evaluate them under the same evaluation setting mentioned above for reconstruction quality. We report inference speed on Nvidia A100 GPU and iPhone 16 Pro Max.
As shown in Tab. 3 and Fig. 1, our models achieve better reconstruction under the same or higher compression ratio. Specifically, with $4\times$ higher compression ratio, our $4\times 16\times 16$ VAE outperforms Cosmos-Tokenizer ($4\times 8\times 8$) across both evaluation sets, achiving notable improvements of +1.76+1.76 PSNR, +0.0333+0.0333 SSIM, 14.22-14.22 rFVD on DAVIS, as well as and +2.27+2.27 PSNR, +0.0174+0.0174 SSIM, 0.96-0.96 rFVD on OpenVid-HD. Furthermore, our $8\times 32\times 32$ VAE significantly outperforms LTX-VAE by +1.89+1.89 PSNR, +0.048+0.048 SSIM, 42.61-42.61 rFVD on DAVIS, along with +0.61+0.61 PSNR, +0.0152+0.0152 SSIM, 0.27-0.27 rFVD on OpenVid-HD. In addition, our model demonstrates superior overall efficiency. Our $8\times 32\times 32$ VAE requires only half the parameters compared to LTX-VAE, and achieves $2.5\times$ speedup on iPhone.
@@ 4.1 Comparison with SOTA autoencoders
Figure 3: AE Qualitative Results. Reconstructions from our H3AE ($8\times 32\times 32$) and other high compression autoencoders: Cosmos-Tokenizer (Agarwal et al., 2025) ($8\times 16\times 16$), LTX-VAE (HaCohen et al., 2024) ($8\times 32\times 32$). We show zoomed-in results to highlight the differences in fidelity and quality. Our method features greater high-frequency detail. GT refers to the ground truth video.
4.2 Ablation Study
We evaluate the effectiveness of the proposed omni-objective training in Sec. 4.2. Compared to the baseline, our approach yields substantial improvements in the image-conditioned I2V setting, while also enhancing the plain T2V setting as a byproduct—effectively providing a “free” gain in reconstruction quality.
We further ablate the proposed latent consistency loss in Sec. 4.2. Results show that it consistently outperforms commonly used auxiliary objectives, such as perceptual and adversarial losses, on both reconstruction accuracy and fidelity metrics, while maintaining greater training stability.
Table 4: Omni-AE for the T2V and I2V tasks. The baseline is trained without image condition while omni-training utilized the image condition with probability ($p=0.5$).
[…]
$8\times 32\times 32$ 29.48 0.8280 0.2627 0.2611 147.1
omni-plain 29.95 0.8367 0.2506 0.2470 129.6
omni-I2V 30.08 0.8384 0.2456 0.2427 128.4
$8\times 64\times 64$ 25.21 0.7284 0.3937 0.4048 737.9
omni-plain 25.97 0.7387 0.3963 0.4041 709.7
omni-I2V 26.38 0.7503 0.3748 0.3879 684.1
Table 5: Training loss comparison. Baseline is $L_{\text{recon}}+\lambda L_{\text{KL}}$. Auxiliary losses and our latent consistency loss are applied to the same base model and trained for 10K iters for comparison.
[…]
Baseline 29.48 0.8280 0.2627 0.2611 147.1
LPIPS 29.07 0.8205 0.224 0.2275 148.9
GAN 29.35 0.8241 0.2631 0.2623 161.0
DWT 29.36 0.8247 0.2651 0.2642 159.3
Ours 29.58 0.8299 0.2541 0.2511 142.2
4.3 Video Generation
We show text-to-video generation results in Tab. 6 and Fig. 4 to demonstrate that our VAE creates a good and highly-compressed latent space for video diffusion models. We construct a 2B DiT aligned with popular video diffusion model settings. Specifically, we use 3D full attention with RoPE and QK-Norm in transformer blocks. All models are first trained on image dataset to learn spatial knowledge for $50K$ iterations, then these model are trained with image-video joint training strategy for $100K$ iterations. The VDM training is done on 128 A100 GPUs for 4-10 days depending on the VAE compression ratio. As shown in Tab. 6, we find that using our H3AE ($8\times 32\times 32$) outperforms Cosmos-Tokenizer ($8\times 16\times 16$) and LTX-VAE in Vbench score. To isolate the impact of the training dataset, we report the scores of both the official LTX model and our finetuned version. Our high-compression ratio autoencoder also enables fast and memory-efficient inference on both GPU and mobile. We additionally report the result of using an extremely high-compression ratio ($8\times 64\times 64$) H3AE, which provides more than a $4\times$ speed-up on the iPhone 16 PM, which demonstrates the great potential of high compression VAEs for efficient video generation.
Table 6: Quantitative Metrics on VBench. Text-to-video generation benchmark on Vbench (Huang et al., 2024) with a 2B DiT denoiser trained using different Autoencoders. Overall Vbench score and selected score are reported. We specifically exhibit aesthetic quality (AQ), imaging quality (IQ), and motion smoothness (MS) to evaluate the performance of denoisers. Latency of DiT is benchmarked by generating 65-frame $512\times 512$ video clips on an NVIDIA A100 80GB GPU and 17-frame $512\times 512$ clips on iPhone 16 Pro Max. Notably the DiT using Cosmos-Tokenizer results in OOM error on iPhone due to memory inefficiency.
@@ 4.3 Video Generation
1LTX-Video DiT finetuned on our datasets.